### v1

# Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance

Abstract
In the age of data-driven medicine, it is paramount to include explainable and ethically managed artificial intelligence in explaining clinical decision support systems to achieve trustworthy and effective patient care. The focus of this paper is on a new architecture of a multi-agent system for clinical decision support that uses modular agents to analyze laboratory results, vital signs, and the clinical context and then integrates these results to drive predictions and validate outcomes. We describe our implementation with the eICU database to run lab-analysis-specific agents, vitals-only interpreters, and contextual reasoners and then run the prediction module and a validation agent. Everything is a transparent implementation of business logic, influenced by the principles of ethical AI governance such as Autonomy, Fairness, and Accountability. It provides visible results that this agent-based framework not only improves on interpretability and accuracy but also on reinforcing trust in AI-assisted decisions in an intensive care setting.
Index Terms:
Clinical Decision Support, Mortality Prediction in the ICU, Deep Learning, Transparency, Large Language Model, AI Agent
I Introduction
AI has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7]. A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], CTs [10], and MRI scans [11], making interpretation faster and more consistent. But it doesn’t stop with images—AI is also being used to make sense of physiological signals, including EEG [12, 13], ECG [14], and data from wearable sensors [15, 16]. These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight—key qualities especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI governance to uphold both safety and accountability.
Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents. Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes. This structure mimics how doctors work—gathering evidence from various sources, weighing context, and forming a unified clinical picture.
By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable and scalable the model is, but also how it upholds fairness and accountability throughout the clinical decision-making process. To test the framework, we draw on the eICU Collaborative Research Database [18], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
II Related Work
II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
Clinical Decision Support Systems (CDSS) have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20]. More recent developments have looked to Clinical Practice Guidelines (CPGs) as a way to enrich Large Language Models (LLMs), boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
Meanwhile, multi-agent approaches to CDSS have gained traction as well. One notable design introduced a case-based reasoning (CBR) framework structured around agents that handle user interaction, task execution, and domain knowledge [20]. Combining MAS with CBR has been shown to improve how efficiently the system learns and adapts to the unique traits of each patient.
II-B eICU Data and Its Applications
The eICU Collaborative Research Database has become a cornerstone for ICU-related studies, offering detailed information from more than 200,000 ICU stays across the United States [21]. With data ranging from vital signs and treatment plans to severity scores, diagnoses, and interventions, it provides a rich foundation for building and evaluating AI models tailored to critical care environments.
The electronic ICU (eICU) itself is a telemedicine-driven model aimed at delivering continuous critical care for high-risk patients [22]. Systems like Philips’ eICU extend ICU expertise to the bedside through the eCareManager platform, which connects hospital networks and generates real-time clinical insights [23].
Deploying eICU systems has led to tangible benefits. At Baptist Health South Florida, for example, implementation was followed by a 23% drop in ICU mortality rates and up to a 25% reduction in patient length of stay [23].
II-C LLM-Based Agents in Healthcare
The rapid advancement of Large Language Models (LLMs) has led to their growing presence in the healthcare space. These models now power a variety of applications—ranging from virtual assistants and personalized health education to symptom checkers and mental health support tools [24]. By streamlining both patient care and administrative tasks, LLM-based agents are poised to reshape how services are delivered.
One notable example is MDAgents, a multi-agent system built on LLMs that tackles complex medical decision-making tasks [25]. Its architecture is designed to reflect how healthcare teams operate in real-world settings, enabling dynamic collaboration among agents. Performance benchmarks show that MDAgents outperforms earlier models across multiple evaluation criteria.
A recent review takes a deep dive into the evolving landscape of LLM-driven agents in medicine [26], covering their technical underpinnings, practical uses, and ongoing challenges. The survey highlights key system elements like planning mechanisms, reasoning strategies, tool integrations, and design configurations. These agents are already being applied in areas such as clinical support, documentation automation, simulation training, and workflow optimization.
II-D Multi-Agent Systems in Healthcare
Multi-Agent Systems (MAS) are gaining attention as a promising way to tackle complex challenges in healthcare. One example applies MAS to pre-hospital emergency response, where agents—such as EMS dispatch centers, ambulances, traffic nodes, and medical providers—collaborate within a distributed decision-making setup [27].
The idea of multi-stage AI agents builds on this by organizing intelligent agents into layers, each handling different parts of perception and reasoning. Many of these layers are now powered by LLMs, allowing for more structured and scalable workflows [28]. This layered setup has shown particular promise in areas like personalized care and remote health services.
Real-world implementations are emerging as well. The LLM-Medical-Agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [29].
II-E Ethical Governance and Explainability in Healthcare AI
As AI continues to find its place in healthcare, the need for ethical oversight and transparent reasoning becomes more urgent. Explainable AI (XAI) plays a key role here by making machine-generated decisions easier for humans to understand—helping both clinicians and patients build confidence in AI-supported care [30]. By shedding light on the logic behind predictions, XAI tackles the long-standing “black-box” issue in conventional AI systems.
Growing concerns around the safety of LLM-based agents have prompted the development of frameworks like GuardAgent [31], which embed policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [32].
Bringing LLMs into electronic health record systems also introduces a range of ethical, legal, and practical questions. These include how to handle consent, maintain oversight, and ensure data governance [33]. A patient-centered approach—with transparency and strong ethical foundations—is essential for protecting vulnerable groups.
To that end, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, wellbeing, and system transparency [34]. Especially in high-risk areas like intensive care units, addressing these governance challenges is key to the responsible deployment of AI [35].
II-F Motivation and Research Gap
While clinical decision support systems, LLM-powered agents [36], and multi-agent frameworks have made notable strides in healthcare, there’s still a considerable gap when it comes to integrating these technologies into real-world ICU settings under clear ethical oversight. Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
Most existing research tends to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified system that reflects how clinicians actually work as a team. This fragmentation underscores the need for a new approach.
In response, we propose a novel agent-based architecture that breaks down the clinical reasoning process into specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout. By embedding ethical AI principles directly into each stage of the pipeline and validating our design using the eICU database, we aim to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
Figure 1: Illustration of the Multi-Agent Framework Design. The system consists of a set of specialized agents, each responsible for processing a specific type of clinical data. The Context Analysis Agent handles unstructured inputs like clinical notes, while the Vitals Analysis Agent focuses on real-time physiological signals, and the Lab Analysis Agent interprets laboratory test results. These distinct streams of information are brought together by the Integration Agent, which fuses multimodal features into a unified representation. Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating length of stay. To support interpretability, the Transparency Agent generates human-readable explanations of model outputs. Finally, the Validation Agent oversees performance assessment by comparing predictions against ground truth data.
III Methods
III-A Dataset and Preprocessing
We based our study on the eICU Collaborative Research Database v2.0 [21], which provides a wide range of de-identified ICU data from over 200,000 admissions across various U.S. hospitals. This dataset combines both structured information—like vital signs and lab results—and unstructured clinical notes, including nursing and physician documentation, offering a comprehensive view of each patient’s stay.
To build our pipeline, we worked directly with a set of key tables from the eICU database, including patient.csv.gz, lab.csv.gz, vitalPeriodic.csv.gz, note.csv.gz, and medication.csv.gz, along with APACHE-related files like apacheApsVar.csv.gz and apachePatientResult.csv.gz. We grouped records by patientunitstayid to keep everything aligned per ICU stay. Incomplete entries—such as patients without enough data in labs, vitals, or notes—were excluded to ensure each case provided a meaningful input for modeling.
The preprocessing wasn’t overly complicated but required care. We cleaned up missing values, sorted events based on when they happened (using time offsets), and trimmed overly long text fields so they’d fit within the token limits of our language model components. The goal here wasn’t just to make the data model-ready, but to keep as much useful clinical context as possible without overwhelming the system.
From the full dataset, we extracted several types of information per patient to construct meaningful input features. For vital signs, we collected the latest ten time-ordered readings per patient to preserve recent physiological trends. In the laboratory domain, we selected the most recent non-redundant values for a set of clinically relevant biomarkers. To capture context from unstructured text, we sampled up to three representative clinical notes per patient, focusing on physician and nursing documentation.
For treatments and medications, we focused on the top 20 items either by uniqueness or frequency across the cohort. Lastly, APACHE scores and predicted outcomes were included as reference metrics for downstream validation and benchmarking.
III-B Multi-Agent Framework Design
To emulate real-world ICU decision-making, we implemented a modular multi-agent architecture consisting of six discrete agents, each responsible for a semantically distinct task. The architecture is shown in Figure 1.
  • Lab Analysis Agent: Receives structured lab data and highlights key abnormalities (e.g., hyperlactatemia, creatinine elevation) with implications on APACHE scoring and patient prognosis.
  • Vitals Analysis Agent: Processes vital signs (e.g., HR, SBP, SpO2, temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
  • Context Analysis Agent: Analyzes unstructured notes, medication usage, and treatment strings to infer diagnoses, risk factors, and progression trajectory.
  • Integration Agent: Aggregates all above agent outputs into a comprehensive, system-by-system clinical assessment. It prioritizes ICU risk factors related to mortality and length of stay (LOS).
  • Prediction Agent: Generates structured outcome predictions (mortality probability and ICU LOS) using integrated findings and APACHE variables. Outputs follow a strict template for automated parsing.
  • Validation Agent: Compares predicted vs. actual ICU outcomes and reflects on the prediction’s accuracy, key contributing variables, and future improvement insights.
Each agent is implemented using Anthropic Claude 3.7 Sonnet and configured via the intelli.flow framework, which allows asynchronous agent orchestration with JSON-structured prompts and logging.
III-C Few-Shot Learning Example Construction
To ground the reasoning of the Prediction Agent, we incorporated two real ICU patients as few-shot exemplars. These examples span different outcomes (e.g., survived vs. expired) and are selected based on APACHE completeness and data richness. Each example includes demographics, APACHE variables, labs, vitals, and actual outcomes. The examples are embedded directly in the prompt using clearly segmented format blocks and used to improve model generalizability.
III-D System Execution and DAG Orchestration
The entire multi-agent pipeline is expressed as a directed acyclic graph (DAG), where tasks are mapped via semantic dependencies. Specifically:
  • lab_analysis, vitals_analysis, and context_analysis feed into integration.
  • integration feeds into prediction.
  • prediction feeds into validation.
Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. Flow state is logged per patient and visualized via networkx-based flow diagrams saved as PNGs. The system supports multi-threaded batch evaluation and error-tolerant retries.
III-E Implementation and Agent API Configuration
Each agent is instantiated as a Claude-based text agent with a predefined mission, API credentials, and output format enforcement. Prompts are customized per task using structured sections (e.g., “KEY ABNORMALITIES”, “APACHE RELEVANT FINDINGS”). Inputs are truncated or summarized to fit within the 10,000-token limit of Claude.
Agent configuration example:
Agent(
  provider="anthropic",
  mission="Analyze lab data for
  abnormalities",
  model_params={"key": ANTHROPIC_API_KEY,
  "model": "claude-3-7-sonnet-20250219"}
All patient data is saved per analysis run, including intermediate and final agent outputs in JSON format.
III-F Ethical AI Governance and Explainability
To ensure safety, fairness, and transparency, the system incorporates several governance mechanisms:
  • Explainability: All agent outputs follow enforced templates, allowing users to trace predictions back to specific findings.
  • Fairness: Demographic data (age, gender, ethnicity) are included in prompts to support bias analysis.
  • Autonomy Boundaries: Agents are not allowed to override downstream agents without structured justification.
  • Auditability: Logs, JSON outputs, and flow diagrams are archived per patient to support reproducibility.
III-G Quantitative the Agent Transparency
The TransparencyMetrics [37] class evaluates the transparency of clinical prediction explanations by analyzing text responses for key transparency features. It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification. The scoring process uses regular expression pattern matching to detect relevant terms in the response content. The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
IV Results
Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error. Each model was executed 10 times, and the box plots represent the distribution of results over these runs. The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models. These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
The experimental results in Table I and Figure 2 indicate that both the multi-agent and single-agent approaches yield remarkably similar performance across all evaluated metrics. Each experiment was conducted ten times per setting, and the reported values represent the average across these ten runs to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 85.54%, while the single-agent model reached 85.52%. Although the multi-agent system shows a slightly higher mean, the difference of 0.02% is minimal and likely not statistically significant. Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 10 runs)
Metric Model Mean Std
Mortality Accuracy (%) Multi-agent 85.543 0.00088
Single-agent 85.520 0.00075
LOS Mean Error (days) Multi-agent 0.74911 0.00107
Single-agent 0.74906 0.00092
LOS Median Error (days) Multi-agent 0.74987 0.00167
Single-agent 0.74882 0.00147
When evaluating the prediction of length of stay (LOS), both the mean and median error metrics show almost no practical difference between the two approaches. The mean LOS error is identical for both models at approximately 0.7491 days, and the median LOS error differs only slightly—0.7499 days for the multi-agent model versus 0.7488 days for the single-agent. Again, the single-agent model demonstrates slightly lower variance, but the magnitude of the difference remains negligible.
TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 10 runs)
Metric Model Mean
Average transparency score (%) Multi-agent 60.01
Single-agent 75.04
Table II presents a comparison of the average transparency scores achieved by the multi-agent and single-agent models across 10 independent runs. The results indicate a clear advantage for the single-agent approach, which attained a significantly higher average transparency score of 75.04%, compared to 60.01% for the multi-agent model. This suggests that the single-agent architecture produces more interpretable or explainable outputs under the current evaluation framework. One possible explanation is that, without explicit inter-agent communication or coordination strategies, the multi-agent model’s internal decision-making process may become more distributed and opaque, reducing its overall transparency. In contrast, the centralized structure of the single-agent model may allow for more consistent and traceable reasoning paths, leading to better transparency metrics.
Overall, the performance of the multi-agent and single-agent models is effectively equivalent in terms of predictive accuracy and error, with only marginal differences that do not clearly favor one approach over the other. Interestingly, in this case, the single-agent model shows slightly better results. This may be due to the absence of explicit interaction mechanisms between agents, which limits the potential benefits of distributed learning. Without proper coordination or communication strategies, the multi-agent system may not fully leverage its structure, making centralized processing by a single agent more effective. This is an insightful finding that highlights how, in the absence of well-designed inter-agent dynamics, simpler single-agent architectures can outperform more complex multi-agent frameworks. Unless there are additional considerations—such as scalability, modularity, or deployment constraints—that specifically favor the multi-agent setup, either model could be considered a valid choice based on the observed outcomes.
V Conclusion
In this study, we compared the performance of single-agent and multi-agent models within a predictive learning framework. Our results show that both approaches achieve comparable levels of predictive accuracy and error, with only marginal differences observed. Interestingly, the single-agent model consistently performed slightly better. We hypothesize that this may be due to the absence of explicit interaction or coordination mechanisms between agents in the current multi-agent setup. Without such mechanisms, the benefits of distributed learning cannot be fully realized, and centralizing all information within a single agent may lead to more effective learning.
This finding highlights an important consideration: multi-agent architectures do not inherently guarantee improved performance unless supported by well-designed inter-agent communication or collaboration strategies. Therefore, unless future system requirements demand advantages such as scalability, modularity, or robustness—areas where multi-agent systems typically excel—the simpler single-agent design may be a more practical and effective choice under current conditions.