### v2

# Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance

Abstract
In the age of data-driven medicine, it is paramount to include explainable and ethically managed artificial intelligence in explaining clinical decision support systems to achieve trustworthy and effective patient care. The focus of this paper is on a new architecture of a multi-agent system for clinical decision support that uses modular agents to analyze laboratory results, vital signs, and the clinical context and then integrates these results to drive predictions and validate outcomes. We describe our implementation with the eICU database to run lab-analysis-specific agents, vitals-only interpreters, and contextual reasoners and then run the prediction module and a validation agent. Everything is a transparent implementation of business logic, influenced by the principles of ethical AI governance such as Autonomy, Fairness, and Accountability. It provides visible results that this agent-based framework not only improves on interpretability and accuracy but also on reinforcing trust in AI-assisted decisions in an intensive care setting.
Index Terms:
Clinical Decision Support, Mortality Prediction in the ICU, Deep Learning, Transparency, Large Language Model, AI Agent
I Introduction
Artificial intelligence (AI) has steadily made its way into many industries, from image recognition and supply chain logistics [1, 2, 3], writing assist [4] to the field of healthcare [5, 6]. Within medicine, it’s proving valuable for sharpening diagnostic precision, supporting treatment planning, and helping clinicians keep a closer eye on patients [7]. A growing body of work has focused on using AI to interpret complex medical visuals like surgical footage [8], X-rays [9], computed tomography (CT) scans [10], and magnetic resonance imaging (MRI) scans [11], making interpretation faster and more consistent. But it doesn’t stop with images. AI is also being used to make sense of physiological signals, including electroencephalography (EEG) [12, 13], electrocardiography (ECG) [14], and data from wearable sensors [15, 16]. These efforts are opening up new possibilities across neurology, psychiatry, and continuous patient monitoring [17]. Altogether, these advancements point to a future where AI supports both visual and signal-based insights, forming the backbone of smarter clinical decision-making tools.
Clinical decision support systems (CDSS) have become a vital part of today’s healthcare settings, offering insights drawn from electronic health records (EHRs) and real-time monitoring tools. Yet, many of the traditional AI models used in these systems fall short when it comes to flexibility, transparency, and oversight key qualities, especially critical in high-risk settings like intensive care units (ICUs). To address these limitations, we introduce a modular multi-agent system (MAS) designed to reflect how clinical teams make decisions, with a built-in emphasis on ethical AI to uphold both explainability and accountability.
Building on progress in agent-based frameworks and the coordinated use of large language models (LLMs), our system breaks down the decision-making pipeline into focused, collaborative agents. Each one is responsible for a different aspect of ICU assessment: from interpreting lab results and tracking vital signs to making context-sensitive judgments based on a patient’s history or co-existing conditions. These individual agents pass their findings to a central integration agent that brings everything together, enabling more comprehensive predictions and cross-validated outcomes. This structure mimics how doctors gathering evidence from various sources, weighing context, and forming a unified clinical picture.
By structuring the system around modular agents and grounding it in ethical oversight, we improve not just how interpretable and scalable the model is, but also how it upholds fairness and accountability throughout the clinical decision-making process. To test the framework, we draw on the eICU Collaborative Research Database [18], showing that our method can deliver well-organized predictions, shed light on key prognostic indicators, and build greater trust in AI-supported medical judgments.
II Related Work
II-A Applications of Clinical Decision Support Systems in Intensive Care Settings
Clinical decision support systems (CDSS) have come a long way, especially in ICU environments where every second counts. Earlier systems typically leaned on rule-based logic or statistical methods to generate recommendations [19, 20]. More recent developments have looked to clinical practice guidelines (CPGs) as a way to enrich LLMs, boosting their ability to offer context-aware treatment advice. Research suggests that LLMs enhanced with CPGs outperform traditional models in delivering more accurate clinical suggestions [19].
Meanwhile, multi-agent approaches to CDSS have gained traction as well. One notable design introduced a case-based reasoning (CBR) framework structured around agents that handle user interaction, task execution, and domain knowledge [20]. Combining MAS with CBR has been shown to improve how efficiently the system learns and adapts to the unique traits of each patient.
II-B eICU Data and Its Applications
The eICU Collaborative Research Database has emerged as a critical resource for intensive care research, gathering comprehensive data from over 200,000 ICU stays across the United States [21]. This extensive collection spans vital parameters—including vital signs, treatment protocols, severity indices, diagnoses, and interventions—serving as a solid groundwork for developing and validating AI models that address the specific challenges of critical care.
The electronic ICU relies on a telemedicine framework to monitor high-risk patients and deliver critical care, even when onsite specialists are not available [22]. For instance, Philips has developed an eICU system that utilizes the eCareManager platform to bring expert ICU support directly to the patient’s bedside. By linking hospital networks and supplying real-time clinical feedback, this approach narrows the gap between off-site experts and immediate patient needs [23].
In practical terms, the rollout of eICU systems has produced measurable improvements in patient outcomes. At Baptist Health South Florida, the introduction of the eICU model was linked to a 23% decrease in ICU mortality and a reduction in the average length of stay by up to 25% [23]. These results not only demonstrate how telemedicine can streamline critical care delivery but also highlight the potential for such systems to transform clinical practice through enhanced real-time decision-making and optimized resource use.
II-C LLM-Based Agents in Healthcare
Large language models (LLMs) have recently become more common in healthcare. These models now assist in multiple areas, including virtual assistants, individualized health education, symptom checking, and mental health support tools [24]. By improving patient interactions and simplifying administrative tasks, LLM-based systems are beginning to influence how healthcare is provided.
One significant example is MDAgents, a multi-agent system using LLMs to manage complex medical decisions [25]. Its design replicates the teamwork observed in actual healthcare environments, enabling effective communication among its agents. Testing has shown that MDAgents performs better than earlier models in various evaluations.
A recent review explored the use of LLM-based agents in medical contexts [26]. The review covered technical foundations, practical applications, and existing challenges. It emphasized components such as planning techniques, reasoning strategies, integration of external tools, and agent architecture. These systems are now employed for tasks like clinical decision support, automatic patient documentation, simulation training, and workflow optimization.
Researchers are now further developing LLM-based agents into multi-agent systems (MAS), where multiple agents interact in a decentralized and collaborative way. This change allows for systems that are more organized and flexible, offering new ways to manage challenging healthcare situations, such as emergency response coordination and personalized patient treatment.
II-D Multi-Agent Systems in Healthcare
Multi-agent systems (MAS) are gaining attention as a promising way to tackle complex challenges in healthcare. One example applies MAS to pre-hospital emergency response, where agents—such as EMS dispatch centers, ambulances, traffic nodes, and medical providers—collaborate within a distributed decision-making setup [27].
The idea of multi-stage AI agents builds on this by organizing intelligent agents into layers, each handling different parts of perception and reasoning. Many of these layers are now powered by LLMs, allowing for more structured and scalable workflows [28]. This layered setup has shown particular promise in areas like personalized care and remote health services.
Real-world implementations are emerging as well. The LLM-medical-agent framework, for instance, demonstrates how MAS can be applied to modular analysis of healthcare data in practical settings [29].
II-E Ethical Governance in Healthcare AI
As AI continues to find its place in healthcare, the need for ethical oversight and transparent reasoning becomes more urgent. Explainable AI (XAI) plays a key role here by making machine-generated decisions easier for humans to understand—helping both clinicians and patients build confidence in AI-supported care [30]. By shedding light on the logic behind predictions, XAI tackles the long-standing “black-box” issue in conventional AI systems.
Growing concerns around the safety of LLM-based agents have prompted the development of frameworks like GuardAgent [31], which embed policy guardrails to ensure compliance with safety and privacy standards—an especially important safeguard in clinical environments [32].
Bringing LLMs into electronic health record systems also introduces a range of ethical, legal, and practical questions. These include how to handle consent, maintain oversight, and ensure data governance [33]. A patient-centered approach—with transparency and strong ethical foundations—is essential for protecting vulnerable groups.
To that end, the World Health Organization (WHO) has outlined ethical guidelines for AI in healthcare, highlighting principles such as human autonomy, wellbeing, and system transparency [34]. Especially in high-risk areas like intensive care units, addressing these governance challenges is key to the responsible deployment of AI [35].
II-F Motivation and Research Gap
While clinical decision support systems, LLM-powered agents and multi-agent frameworks [36], have made notable strides in other fields, there’s still a considerable gap when it comes to integrating these technologies into healthcare, especially for real-world ICU settings under clear ethical oversight. Many current solutions fall short in modularity, lack transparency, or aren’t built with the kind of structured, inter-agent communication needed to reflect the fast-paced, interdisciplinary nature of intensive care.
Most existing research tends to focus on isolated tasks—like interpreting lab results, monitoring vital signs, or reasoning based on medical history—but few bring these components together into a unified system that reflects how clinicians actually work as a team. This fragmentation underscores the need for a new approach.
In response, we propose a novel agent-based architecture that breaks down the clinical reasoning process into specialized, collaborative agents—each designed to handle a distinct aspect of care while maintaining accountability and interpretability throughout. By embedding ethical AI principles directly into each stage of the pipeline and validating our design using the eICU database, we aim to bridge both the technical and ethical gaps in deploying trustworthy AI for high-stakes decision-making in critical care.
Figure 1: Illustration of the Multi-Agent Framework Design. The system consists of a set of specialized agents, each responsible for processing a specific type of clinical data. The Context Analysis Agent handles unstructured inputs like clinical notes, while the Vitals Analysis Agent focuses on real-time physiological signals, and the Lab Analysis Agent interprets laboratory test results. These distinct streams of information are brought together by the Integration Agent, which fuses multimodal features into a unified representation. Based on this, the Prediction Agent carries out key forecasting tasks—such as predicting ICU mortality or estimating length of stay. To support interpretability, the Transparency Agent generates human-readable explanations of model outputs. Finally, the Validation Agent oversees performance assessment by comparing predictions against ground truth data.
III Methods
III-A Dataset and Preprocessing
This study used the eICU Collaborative Research Database v2.0 [21], which compiles anonymized ICU records from over 200,000 patient admissions in diverse U.S. hospitals. The database encompasses structured details (e.g., vital signs and lab results) and unstructured clinical notes contributed by nurses and physicians, giving us a thorough view of patient care.
We concentrated on several key eICU files: patient.csv, lab.csv, vitalPeriodic.csv, note.csv, and medication.csv, along with APACHE-related data files [37] (apacheApsVar.csv and apachePatientResult.csv). To align each patient’s information, records were grouped according to patient-unit-stayid. If any essential data were missing—such as vital signs, lab values, or clinical notes—we removed those entries to maintain reliability.
We addressed the data gaps, ordered events based on their timestamps, and shortened lengthy text fields to meet language model guidelines. Then, we sampled 150 patients for the study: 76 mortality patients and 74 survived patients.
Next, we extracted specific features from each patient’s record set. We captured the ten most recent vital sign readings to reveal each patient’s current physiological state. We also selected the latest distinct lab biomarkers deemed clinically relevant. When dealing with unstructured clinical documentation, we included up to three notes for every patient, focusing largely on entries written by physicians and nurses.
Our analysis also tracked the top 20 medications and treatments, identifying them by frequency or uniqueness within the overall dataset. Finally, we incorporated APACHE scores and predictions as reference points to aid in validating and evaluating our modeling outcomes.
III-B Multi-Agent Framework Design
To emulate real-world ICU decision-making, we implemented a modular multi-agent architecture consisting of six discrete agents, each responsible for a semantically distinct task. The architecture is shown in Figure 1.
  • Lab Analysis Agent: Receives structured lab data and highlights key abnormalities (e.g., hyperlactatemia, creatinine elevation) with implications on APACHE scoring and patient prognosis.
  • Vitals Analysis Agent: Processes vital signs (e.g., heart rate (HR), systolic blood pressure (SBP), peripheral capillary oxygen saturation ($SpO_{2}$), temperature) and evaluates physiological stability, respiratory function, and cardiovascular performance.
  • Context Analysis Agent: Analyzes unstructured notes, medication usage, and treatment strings to infer diagnoses, risk factors, and progression trajectory.
  • Integration Agent: Aggregates all above agent outputs into a comprehensive, system-by-system clinical assessment. It prioritizes ICU risk factors related to mortality and length of stay (LOS).
  • Prediction Agent: Generates structured outcome predictions (mortality probability and ICU LOS) using integrated findings and APACHE variables. Outputs follow a strict template for automated parsing.
  • Validation Agent: Compares predicted vs. actual ICU outcomes and reflects on the prediction’s accuracy, key contributing variables, and future improvement insights.
To maximize information flow between agents, we implemented a shared memory architecture that allows any agent to access inputs and outputs from previous pipeline stages. This approach mitigates the risk of information loss between modules while maintaining the semantic separation of responsibilities. While this shared memory design has proven effective at improving predictive performance by ensuring that no critical clinical insights are lost during the analysis process, we recognize that the multi-agent architecture introduces additional complexity that can impact transparency. Our current implementation focuses on performance optimization, with ongoing work to enhance the explainability mechanisms across the agent communication pathways.
Each agent is implemented using OpenAI GPT-4o [38] and configured via the intelli framework [39], which allows asynchronous agent orchestration with JSON-structured prompts and logging.
III-C Few-Shot Learning Example Construction
To ground the reasoning of the Prediction Agent, we incorporated two real ICU patients as few-shot exemplars. These examples span different outcomes (e.g., survived vs. expired) and are selected based on APACHE completeness and data richness. Each example includes demographics, APACHE variables, labs, vitals, and actual outcomes. The examples are embedded directly in the prompt using clearly segmented format blocks and used to improve model generalizability.
III-D System Execution and DAG Orchestration
The entire multi-agent pipeline is expressed as a directed acyclic graph (DAG), where tasks are mapped via semantic dependencies. Specifically:
  • lab_analysis, vitals_analysis, and context_analysis feed into integration.
  • integration feeds into prediction.
  • prediction feeds into validation.
Execution is managed asynchronously using Python’s asyncio to allow concurrent LLM calls and reduce latency. The system supports multi-threaded batch evaluation and error-tolerant retries.
Figure 2: Comparison of performance between the Multi-agent and Single-agent frameworks across three evaluation metrics: Mortality Prediction Accuracy, Length of Stay (LOS) Mean Error, and LOS Median Error. Each model was executed 8 times, and the box plots represent the distribution of results over these runs. The Multi-agent framework shows slightly higher mean mortality accuracy with slightly more variance, while LOS-related errors are nearly identical between the two models. These results indicate comparable predictive performance, with minor differences in consistency and central tendency.
III-E Implementation Details
Each agent is instantiated as a generative pre-trained transformer-based (GPT-based) [40] text agent with a predefined mission, API credentials, and output format enforcement. Prompts are customized per task using structured sections (e.g., “KEY ABNORMALITIES”, “APACHE RELEVANT FINDINGS”). Inputs are truncated or summarized to fit within the 10,000-token limit of GPT.
Agent configuration example:
Agent(
  provider="openai",
  mission="Analyze lab data for
  abnormalities",
  model_params={"key": OPENAI_API_KEY,
  "model": "gpt-4o"}
All patient data is saved per analysis run, including intermediate and final agent outputs in JSON format.
III-F Ethical AI and Explainability
To ensure safety, fairness, and transparency, the system incorporates several governance mechanisms:
  • Explainability: All agent outputs follow enforced templates, allowing users to trace predictions back to specific findings.
  • Fairness: Demographic data (age, gender, ethnicity) are included in prompts to support bias analysis.
  • Autonomy Boundaries: Agents are not allowed to override downstream agents without structured justification.
  • Auditability: Logs, JSON outputs, and flow diagrams are archived per patient to support reproducibility.
III-G Quantitative Evaluation of Agent Transparency
Here we built the transparency assessment module [41] to evaluate the transparency of clinical prediction explanations by analyzing text responses for key transparency features. It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification. The scoring process uses regular expression pattern matching to detect relevant terms in the response content. The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
IV Results
The experimental results in Table I and Figure 2 demonstrate that the multi-agent system consistently outperforms the single-agent approach across all evaluated metrics over 150 unique patients. Each experiment was conducted across eight runs with approximately 150 patients per run, and the reported values represent the average performance to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 59%, while the single-agent model reached only 56%. This 3 percentage point improvement is consistent across multiple runs and represents enhancement in predictive capability in the high-stakes ICU environment. Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
TABLE I: Comparison of Multi-agent and Single-agent Models (Average over 8 runs)
Metric Model Mean
Mortality Prediction Accuracy (%) Multi-agent 59.00
Single-agent 56.00
LOS Mean Error (days) Multi-agent 4.37
Single-agent 5.82
Mean Squared Error (days²) Multi-agent 35.49
Single-agent 48.13
Root Mean Squared Error (days) Multi-agent 5.95
Single-agent 6.94
Our analysis indicates an enhancement in predicting Length of Stay (LOS) when using the multi-agent approach. In our study, the average prediction error drops to 4.37 days under the multi-agent strategy, compared to 5.82 days observed with the single-agent method—an improvement of roughly 25% in accuracy. This gain is important, given its direct influence on how ICU resources are allocated and care is planned. In addition, the multi-agent system registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
A particularly aspect of our findings is how the multi-agent approach manages to reduce the LOS prediction error. A mean error of 4.37 days, as opposed to 5.82 days from the single-agent method, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care. Moreover, the improved metrics—lower mean squared error (35.49 instead of 48.13) and root mean squared error (5.95 compared to 6.94)—further confirm that this system not only enhances accuracy but also offers more stable predictions across various patient groups.
Figure 3: Comparison of Multi-agent and Single-agent frameworks across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom).
Figure 3 reveals a clear analysis across the eight test runs we conducted. Looking at the top graph, you can see the Multi-agent system (blue line) consistently outperformed the Single-agent approach (green line) in predicting mortality. The Multi-agent accuracy ranges from about 56% to nearly 60%, while the Single-agent stays between 55%–57%. What’s notable not just that it performed better, but that this advantage held steady across every single run.
The LOS prediction results in the bottom graph show more improvement for multi-agent model. The blue line stays below the green throughout all runs, with errors around 4–4.7 days compared to the Single-agent’s 5.7–6 days. That gap - somewhere around a day and a half - might not sound huge until you consider what it means for real patients and hospital planning. Most systems like this show more ups and downs, but here the multi-agent setup maintained its edge consistently.
TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 8 runs)
Metric Model Mean
Average transparency score (%) Multi-agent 85.50
Single-agent 86.21
Table II provides a side-by-side comparison of average transparency scores from eight independent runs. The data shows that the single-agent approach scores an average of 86.21% in contrast to 85.50% for the multi-agent model. This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the single-agent design being marginally more interpretable. Despite the distributed nature of the multi-agent system, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
Overall, both the multi-agent and single-agent models perform similarly in terms of predictive accuracy and error, with only slight differences that do not strongly favor one over the other. Interestingly, the single-agent model shows a marginal advantage, which might be linked to its avoidance of the complexities that come with managing interactions among multiple agents. Without effective coordination, a multi-agent system may not fully benefit from its distributed learning structure, making centralized processing by a single agent a more effective option under these circumstances. This finding highlights that, in the absence of well-optimized inter-agent dynamics, the simpler single-agent architecture can sometimes outperform a more complex system. Therefore, unless additional factors—such as scalability or deployment constraints—demand a multi-agent setup, both models are viable choices based on the results observed.
V Conclusion
In our recent analysis, we found that the multi-agent system delivers more accurate predictions than the single-agent model. For example, it not only improves mortality prediction rates and shortens the forecasted length of stay—but it does so with comparable interpretability. Specifically, the single-agent model scores 86.21% for transparency, whereas the multi-agent system comes in at a very close 85.50%. This minimal difference suggests that our multi-agent architecture effectively maintains transparency despite the inherent complexity of coordinating decisions among several specialized agents.
This balance between performance and understandability has serious implications in the clinical arena. In scenarios where prediction accuracy is essential—such as in critical care decisions—the multi-agent approach clearly offers benefits. Conversely, when it is crucial to explain the reasoning behind predictions to build clinician trust, the higher transparency of the single-agent model is more attractive.
Looking ahead, our research will concentrate on boosting the clarity of the multi-agent framework without compromising its predictive strengths. By enhancing inter-agent communication protocols and incorporating more advanced explanation mechanisms, we aim to develop a system that marries the superior predictive power of multi-agent architectures with the interpretability required for safe and effective use in critical care.